Sunland Company has a new product going on the market next year. The following data are projections for production and sales: Variable costs $125000 Fixed costs $450000 ROI 14% Investment $1400000 Sales 100000 units What is the target selling price per unit?
Answer #1

ROI=Net operating income/Value of investment

Net operating income=(1400000*14%)=$196000

Target contribution margin=Fixed cost+Net operating income

=196000+450000=$646000

Contribution margin=Sales-Variable cost  

Hence target sales=646000+125000=$771000

Hence target selling price=771000/100000

=$7.71 per unit